A US manufacturing firm borrows $1,500,000 at LIBOR plus a margin of 1.25 percent per annum...

  1. A US manufacturing firm borrows $1,500,000 at LIBOR plus a margin of 1.25 percent per annum on a six-month rollover basis from a UK based bank. If six-month LIBOR is 4 1⁄2 percent over the first six-month interval and 5 3/8 percent over the second six- month interval, how much will the US manufacturing firm pay in interest over the first year of its Eurodollar loan?

  2. Would the US manufacturing firm have been able to obtain cheaper funding had it issued corporate bonds? Which information would you need to be able to answer the question accurately?
Answer #1

Borrowings = $ 1500000, Interest Rate = (LIBOR + 1.25 %) per annum

LIBOR for first 6 months = 4.5 % per annum and LIBOR for second 6 months = 5.375 % per annum

Interest Rate for first 6 months = (4.5 + 1.25) / 2 = 2.875 % and Interest Rate for second 6 months = (5.375 + 1.25) / 2 = 3.3125 %

Interest Expense over the first year = 0.02875 x 1500000 + 0.033125 x 1500000 = $ 92812.5

In order to understand if a bond issuance would have been a cheaper method of financing or not, one would need to know the prevailing interest rate at the time of the bond's individual periodic cash flows. If the interest rates prevailing during the bond tenure is lower than the LIBOR + 1.25 % premium interest rate, the cost of funding in case of a bond issuance will be lower. However, if the bond's interest rate is greater than the variable rate loan, the bond issuance will, in fact, be a costlier method of financing.